# Bistability of travelling waves and wave-pinning states in a mass-conserved reaction-diffusion system:

### v3

@@ 5 Solutions and bifurcations in the nonlinear regime
Linear stability analysis reflects the spatial separation of scales associated with the primary long wavelength (LW) and finite wavenumber Hopf (WB) instabilities. Therefore, in what follows, we find it useful to separate the analysis and investigate solutions in large domains ($L\gg\lambda$) and in moderate domains ($L\sim{\mathcal{O}}(m\lambda$)) along the slice depicted in Figure 4, where $\lambda$ is the critical wavelength of the WB in the codimension-2 LW/WB bifurcation at $(s,b)=(s_{c}^{t},b_{c})$ and $m$ is some small positive integer. The analysis is performed using the numerical continuation package AUTO [25]. For presentation purposes, we plot the bifurcation diagrams with a combination of the standard Sobolev $H^{1}$ norm
@@ 5.1 Coherent solutions in large domains, $L\gg\mathcal{O}(\lambda)$
Figure 8: (a) Bifurcation diagrams computed via numerical continuation as a function of $s$, showing excitable pulses (EPs), travelling fronts (TFs), and travelling waves (TWs), for $M=4.5$ and $L=1000$ near the T-point (square) shown in Fig. 6. The wavelength of the TWs equals that of the oscillations seen in the periodic-$Q_{1}^{\ast}$-$Q_{0}^{\ast}$ connections. The solution branches are projected onto the propagation speed $c$. All solutions in this diagram are linearly unstable. The critical value $s_{T}\approx 1.804$ represents the location of the T-point involving EP solutions. The inset in (a) shows the region where periodic-$Q_{1}^{\ast}$-$Q_{0}^{\ast}$ are observed but is instead projected with the Sobolev norm Eq. 23. Here, we see a snaking behaviour. (b) Selected solution profiles corresponding to the points marked in (a) zoomed in to the spatial oscillations at the left boundary. Each solution shown is a periodic-$Q_{1}^{\ast}$-$Q_{0}^{\ast}$ cycle (see the hollow square and pentagon in Fig. 7b for the full profiles). Here we see that at each fold the spatial oscillations lose half a period. On the circle profile, we overlay the TW solution with five wavelengths showing that the spatial oscillations are travelling waves when isolated on a periodic domain. Note that the TW and TF branches should intersect at the circle but do not because of numerical inaccuracies. Other parameter values as in Table 1 with $b=0.00067$.
@@
For completeness, we also indicate that in the regime $s^{b}_{LW}<s<s_{WB}$, there is bistability between the HSSs $Q_{0}^{\ast}$ and $Q_{2}^{\ast}$ (see Figure 6), which gives rise to TFs. Below the WB instability, $s_{WB}$, fronts connecting $Q_{0}^{\ast}$ and $Q_{2}^{\ast}$ exhibit monotonic bi-asymptotic spatial decay and are linearly stable, whereas for $s>s_{WB}$, the fronts are unstable while exhibiting spatially decaying oscillations about $Q_{2}^{\ast}$. Above $s_{WB}$, these oscillations extend, forming a heteroclinic connection between a spatially periodic orbit and $Q_{0}^{\ast}$. Note that TFs between periodic orbits and homogeneous steady states have been observed in reaction-diffusion systems via numerical integration on a finite domain [77, 78]. Additional increase in $s$ leads to a snaking behavior (see Fig. 8), in the vicinity of $s=s_{T}$, where the solutions resemble double-heteroclinic states (in the comoving frame): periodic-$Q_{1}^{\ast}$-$Q_{0}^{\ast}$ (see Fig. 7c). The periodic orbits in the snaking region are TWs and in Fig. 8, we also superimpose a TW branch with a wavelength matching that of the periodic orbits. Along the snaking branches, the propagation speeds of the solutions oscillate as the periodic orbit is pushed out of the (finite) domain. The solutions lose one period of the left oscillations for every complete cycle from circle to hollow square back to circle in Fig. 8. The inset in Fig. 8 shows a snaking structure observed using the Sobolev norm projection Eq. 23. At the initial fold (hollow square) there are twelve full periods of the oscillation (with this domain length $L=1000$). At the next fold (circle), half a period is lost, and at the next (inverted triangle), there are only eleven wavelengths. This process continues until only half a period remains (hollow pentagon). In the infinite domain, the reverse process of adding more periods at each fold (i.e., travelling up the branch) would occur indefinitely. Since TFs are not spatially reversible solutions, only one period (not two) is added after passing two saddle-nodes, bearing a similarity to [52]. We demonstrate two profiles in Figure 7c (see hollow square and pentagon), and in Figure 7d, we show the respective phase portrait projections.
@@
Figure 9: Bifurcation diagrams computed via numerical continuation as a function of $s$, showing excitable pulses (EPs), wave-pinning (WP) solutions, homogeneous steady states (HSSs) $Q^{\ast}_{0,1,2}$, and the stationary pulses (SP) that emerge from the long wavelength (LW) bifurcations for $M=2$ and $L=1000\gg\lambda\approx 3.09$ (where $\lambda$ is the wavelength leading to instability of $Q_{2}^{\ast}$). The solution branches are projected with Eq. 23 in (a), the maximum value of $F$ in (b), and the speed $c$ in (c); solid (dashed) lines denote linearly stable (unstable) solutions. The insets in (a) show the locations of several folds along the EP branch that are invisible otherwise, and schematics of the spatial eigenvalues $\zeta$ of $Q_{2}^{\ast}$ at ($s=s_{c}^{t}$) and near (pentagon) the LW onset along the upper SP branch. The square denotes two zero eigenvalues due to mass conservation and spatial reversibility. This also holds for the LW of $Q_{0}^{\ast}$. The right inset in (a) shows the eigenvalue configuration of $Q_{0}^{\ast}$ at a stable EP (i.e., with respective speed $c$, inverted hollow triangle), where the circle shows the zero eigenvalue from mass conservation. The inset in (b) shows the eigenvalue configuration for EPs after the parity-breaking bifurcation. The insets in (c) show solutions at the T-points zoomed into the peaks and the separation of the two T-points. The critical values $(s_{c}^{t},s_{LW},s_{HB},s_{T},s_{P},s_{PD})=(0.409,0.548,0.565,0.602,0.684,0.978)$ denote the codimension-2 LW/finite wavenumber Hopf bifurcation (WB), the LW of $Q_{0}^{\ast}$, the homogeneous Hopf (HB), the T-point bifurcation (red square), the parity-breaking bifurcation of SPs (black triangle), and period-doubling (red circle) bifurcations of EPs. The other symbols (pentagon, diamond, etc.) along the branches represent selected locations of solutions given in Fig. 11. The hollow pentagon denotes the location of the HB (see Fig. 10). Other parameter values as in Table 1 with $b=b_{c}\approx 0.067$.
@@
Figure 10: (a) Bifurcation diagram computed via numerical continuation as a function of $s$, showing excitable pulses (EPs), homogeneous steady states $Q^{\ast}_{0,1,2}$, and the long wavelength oscillatory (LWO) solutions that emerge from the homogeneous Hopf bifurcation (HB) for $M=2$ and $L=2000$. The solution branches are projected with Equation 23. Solid (dashed) lines denote linearly stable (unstable) solutions. The left inset shows that the propagation speed $c$ of the EPs tends to zero as it approaches the saddle-node $SN_{Q_{1/2}^{\ast}}$. The right insets show schematics of the spatial eigenvalues $\zeta$ at the onset near the saddle-node $SN_{Q_{1/2}^{\ast}}$ and shortly after where $Q_{1}^{\ast}$ is the background state along the crimson red EP branch. The square denotes the two zero eigenvalues due to mass conservation and spatial reversibility when $c=0$ and the circle shows the persistent zero eigenvalue from mass conservation once spatial reversibility is lost. The critical values $(s_{HB},s_{SN}^{1/2},s_{T})=(0.565,0.601,0.602)$ represent the homogeneous Hopf (HB), saddle-node of $Q_{1/2}^{\ast}$, and the T-point, respectively. (b) Real and (c) magnitude of the imaginary parts of the spatial eigenvalues $\zeta$ of the background state $Q_{2}^{\ast}$ along the red branch of EP solutions given between the T-points (TP, square) and the far-left fold (SN, diamond). Here, $\beta$ denotes the arc length along the red EP branch. The sixth eigenvalue is not shown but is negative, with $|\zeta|\sim\mathcal{O}(100)$. The dashed line in (b) denotes the magnitude of the real part of the negative real part complex conjugate pair (orange). The starting TP is given by red square in Fig. 9 and the ending TP by the bordered square. The other symbols (pentagon, diamond, etc.) along the branches represent selected locations of solutions given in Fig. 11c. Other parameter values as in Table 1 with $b=b_{c}\approx 0.067$.
@@
Figure 11: Solution profiles at selected locations along the (a) EP, (b) SP, and (c) WP branches in Fig. 9, and the (d) EP and (e) LWO branches in Fig. 10, where for visibility, the whole domain length of $L=1000$ is not always shown. Here $(s_{PD},s_{P},s_{T})\approx(0.978,0.684,0.602)$ are defined as in Fig. 9, $s_{SN}^{1/2}=0.601$ as in Fig. 10, and $(c_{T_{1}},c_{T_{2}})\approx(0.260,0.222)$ denote the propagation speeds at the two T-points. Selected phase space projections are shown for EPs in (f) and WP solutions in (g), where in (g) only the homogeneous steady state $Q_{2}^{\ast}$ exists.
@@
where the full system for which the computation is done, is being described in Equation 42. The inset in Fig. 9b shows the spatial eigenvalue configuration about $Q_{0}^{\ast}$ of the EP branch shortly after the parity-breaking bifurcation but before the onset of saddle-focus connections. Here we see that the zero eigenvalue from spatial reversibility becomes negative but initially all eigenvalues are real. Afterwards, the solutions stabilize after the first fold and destabilize at the subsequent left fold (hollow square). The inset in Fig. 9a shows the spatial eigenvalue configuration about $Q_{0}^{\ast}$ of the EP branch (inverted hollow triangle), where we see that the trailing edge has become oscillatory. After the EPs destabilize, the branch folds again before reaching a T-point bifurcation (square) at $s=s_{T}$ (see Fig. 11a). Here, the solutions form a homoclinic cycle with $Q_{0}^{\ast}$ at the saddle-node, i.e., the solutions briefly plateau at $Q_{2}^{\ast}=Q_{1}^{\ast}$ (see square in Fig. 11c). After passing the T-point, the spatially oscillatory tail becomes a leading edge decaying to $Q_{2}^{\ast}$, as shown by the profile at the left fold (diamond), see also the phase portrait projection in Fig. 11c. This switch is possible by additional annihilation of real eigenvalues and formation of a double focus before reaching the left fold. In Fig. 10b,c, we show the spatial eigenvalues, $\zeta$, corresponding to the above transitions. Here we see that the solutions are saddle-focus connections towards the T-points (squares) and focus-focus connections around the far-left fold (diamond). At the HB of the HSS $Q_{2}^{\ast}$, the magnitudes of the real parts of the leading eigenvalues swap and there is a brief regime where saddle-focus Shil-nikov homoclinics emerge [50]. Then, the branch passes the T-point again, where an additional peak appears out of the leading edge oscillations. The branch associated with a bounded two-peak state goes through two more folds (hollow diamond and circle), after which the peaks start to separate until reaching equidistance, and thus, the branch terminates in a period-doubling bifurcation at $s=s_{PD}$ (a consequence of finite size domain). We omit the EP branch with wavelength $L=500$ as the branch is effectively the same as with wavelength $L=1000$. Note that similar bifurcation structures related to peak doubling near T-point bifurcations have also been observed in an RD model but in the absence of mass conservation [47].
@@
Next, we address the steady solution branch that bifurcates from $s=s_{c}^{t}$. As with high mass ($M=4.5$), the solutions emerge subcritically from $Q_{2}^{\ast}$ and $Q_{0}^{\ast}$ but these onsets can be instabilities when $M=2$, as shown in Figures 4 and 9. From both onsets and with both masses, the bifurcating solutions are spatially localized [9, 69] stationary pulses (SP) that rapidly grow in their amplitudes toward the respective folds [16]. From $Q_{2}^{\ast}$, these localizations are holes while from $Q_{0}^{\ast}$, these are peaks and the onset is close to $SN_{Q^{\ast}_{0/1}}$ (this persists for other values of $b$, see also Fig. 3a). Note that similar behaviour is observed at saddle-nodes in dissipative systems without mass conservation [18, 70, 97]. In the insets of Fig. 9a, we show that the spatial eigenvalue configuration about $Q_{2}^{\ast}$ and $Q_{0}^{\ast}$ follows the typical configuration of eigenvalues required for localization, whereby a pair of double zero eigenvalues split into two reals, coexisting with an additional positive/negative real pair (i.e., intersection of two two-dimensional manifolds [46]). After the folds, the hole and the peak widen forming WP solutions, where $s$ controls the width. Hence, in between the two folds, the solutions form a homoclinic-like cycle, resembling two locked fronts. However, both plateaus are not HSSs of the system, as shown for mass-conserved systems in [87] and seen by the profile and phase-portrait projection marked by the hollow pentagon in Fig. 11b,d. Stability computations show that the WP solutions are linearly stable from the left fold and up to $s\approx 0.718$, see Figure 9.
@@
Figure 12: Bifurcation diagrams computed via numerical continuation as a function of $s$, showing the homogeneous steady states $Q^{\ast}_{0,1,2}$, and travelling waves (TWs) and wave-pinning (WP) solutions with wavelengths $\lambda$ and $2\lambda$ for $M=2$ and $L=2\lambda\approx 2\cdot 3.09$ (where $\lambda$ is the critical wavelength at the instability onset of $Q_{2}^{\ast}$). The solution branches are projected with the Sobolev norm Eq. 23 in (a) and the propagation speed $c$ in (b); solid (dashed) lines denote linearly stable (unstable) solutions. The notation 2TWλ denotes two spatial copies of travelling waves with wavelength $\lambda$. Here, TWs and WP solutions form a bistable region with wavelengths $\lambda$ and $2\lambda$, respectively. The critical values $(s_{c}^{t},s_{P}^{\lambda},s_{P}^{2\lambda},s_{PD}^{\lambda})=(0.409,0.593,0.664,0.760)$ denote the codimension-2 long wavelength (LW)/finite wavenumber Hopf (WB) onset, the parity-breaking bifurcations (black triangles) at wavelengths $\lambda$ and $2\lambda$, respectively, and the period-doubling (hollow pentagon) bifurcation of $\lambda$. The symbols (circle, diamond, etc.) represent selected locations of solution profiles given in Fig. 13. Other parameter values as in Table 1 with $b=b_{c}\approx 0.067$.
@@
Figure 14: (a) Bifurcation diagram computed via numerical continuation as a function of $s$, showing the homogeneous steady states $Q^{\ast}_{0,1,2}$, and travelling waves (TWs) and wave-pinning (WP) solution branches with wavelengths $3\lambda/2$ and $3\lambda$ for $M=2$ and $L=3\lambda\approx 3\cdot 3.09$ (where $\lambda$ is the critical wavelength at the instability onset of $Q_{2}^{\ast}$). Here we only show projections with the Sobolev norm Eq. 23; solid (dashed) lines denote linearly stable (unstable) solutions. The notation 2TW3λ/2 denotes two spatial copies of TWs with wavelength $3\lambda/2$, where the onsets of each one of the solution branches are at the respective parity-breaking bifurcations. The critical values are $(s_{c}^{t},s_{P_{l}}^{3\lambda/2},s_{P_{r}}^{3\lambda/2},s_{P}^{3\lambda},s_{H}^{3\lambda/2},s_{PD}^{3\lambda/2})=(0.409,0.449,0.629,0.659,0.604,0.821)$, for the codimension-2 long wavelength (LW)/finite wavenumber Hopf (WB) bifurcation, the parity-breaking bifurcation at wavelengths $3\lambda/2$ on the left and right, and $3\lambda$, respectively, and the Hopf (H) and period-doubling bifurcations of TW3λ/2. (b-e) Selected solution profiles for (b) 2WP3λ/2 (c) 2TW3λ/2 (d) WP and (e) TW at corresponding points marked in (a). Other parameter values as in Table 1 with $b=b_{c}\approx 0.067$.

### v4

@@ 5 Solutions and bifurcations in the nonlinear regime
Linear stability analysis reflects the spatial separation of scales associated with the primary long wavelength (LW) and finite wavenumber Hopf (WB) instabilities. Therefore, in what follows, we find it useful to separate the analysis and investigate solutions in large domains ($L\gg\lambda$) and in moderate domains $L\sim{\mathcal{O}}(m\lambda$)] along the slice depicted in Figure 4, where $\lambda$ is the critical wavelength of the WB in the codimension-2 LW/WB bifurcation at $(s,b)=(s_{c}^{t},b_{c})$ and $m$ is some small positive integer. The analysis is performed using the numerical continuation package AUTO [25]. For presentation purposes, we plot the bifurcation diagrams with a combination of the standard Sobolev $H^{1}$ norm
@@ 5.1 Coherent solutions in large domains, $L\gg\mathcal{O}(\lambda)$
Figure 8: (a) Bifurcation diagrams computed via numerical continuation as a function of $s$, showing excitable pulses (EPs), travelling fronts (TFs), and travelling waves (TWs), for $M=4.5$ and $L=1000$ near the T-point (square) shown in Fig. 6. The wavelength of the TWs equals that of the oscillations seen in the periodic-$Q_{1}^{\ast}$-$Q_{0}^{\ast}$ connections. The solution branches are projected onto the propagation speed $c$; solid (dashed) lines denote linearly stable (unstable) solutions. The critical value $s_{T}\approx 1.804$ represents the location of the T-point involving EP solutions. The inset in (a) shows the region where periodic-$Q_{1}^{\ast}$-$Q_{0}^{\ast}$ are observed but is instead projected with the Sobolev norm Eq. 25. Here, we see a snaking behaviour. (b) Selected solution profiles corresponding to the points marked in (a) zoomed in to the spatial oscillations at the left boundary. Each solution shown is a periodic-$Q_{1}^{\ast}$-$Q_{0}^{\ast}$ cycle (see the hollow square and pentagon in Fig. 7c for the full profiles). Here we see that at each fold the spatial oscillations lose half a period. On the circle profile, we overlay the TW solution with five wavelengths showing that the spatial oscillations are travelling waves when isolated on a periodic domain. Note that the TW and TF branches should intersect at the circle but do not because of numerical inaccuracies. Other parameter values as in Table 1 with $b=0.00067$.
@@
For completeness, we also indicate that in the regime $s^{b}_{LW}<s<s_{WB}$, there is bistability between the HSSs $Q_{0}^{\ast}$ and $Q_{2}^{\ast}$ (see Figure 6), which gives rise to TFs. Below the WB instability, $s_{WB}$, fronts connecting $Q_{0}^{\ast}$ and $Q_{2}^{\ast}$ exhibit monotonic bi-asymptotic spatial decay and are linearly stable, whereas for $s>s_{WB}$, the fronts are unstable while exhibiting spatially decaying oscillations about $Q_{2}^{\ast}$. Above $s_{WB}$, these oscillations extend, forming a heteroclinic connection between a spatially periodic orbit and $Q_{0}^{\ast}$. Note that TFs between periodic orbits and homogeneous steady states have been observed in reaction-diffusion systems via numerical integration on a finite domain [78, 79]. Additional increase in $s$ leads to a snaking behavior (see Fig. 8), in the vicinity of $s=s_{T}$, where the solutions resemble double-heteroclinic states (in the comoving frame): periodic-$Q_{1}^{\ast}$-$Q_{0}^{\ast}$ (see Fig. 7c). The periodic orbits in the snaking region are TWs and in Fig. 8, we also superimpose a TW branch with a wavelength matching that of the periodic orbits. Along the snaking branches, the propagation speeds of the solutions oscillate as the periodic orbit is pushed out of the (finite) domain. The solutions lose one period of the left oscillations for every complete cycle from circle to hollow square back to circle in Fig. 8. The inset in Fig. 8 shows a snaking structure observed using the Sobolev norm projection Eq. 25. At the initial fold (hollow square) there are twelve full periods of the oscillation (with this domain length $L=1000$). At the next fold (circle), half a period is lost, and at the next (inverted triangle), there are only eleven wavelengths. This process continues until only half a period remains (hollow pentagon). In the infinite domain, the reverse process of adding more periods at each fold (i.e., travelling up the branch) would occur indefinitely. Since TFs are not spatially reversible solutions, only one period (not two) is added after passing two saddle-nodes, bearing a similarity to [71, 52]. We demonstrate two profiles in Figure 7c (see hollow square and pentagon), and in Figure 7d, we show the respective phase portrait projections.
@@
Figure 9: Bifurcation diagrams computed via numerical continuation as a function of $s$, showing excitable pulses (EPs), wave-pinning (WP) solutions, homogeneous steady states (HSSs) $Q^{\ast}_{0,1,2}$, and the stationary pulses (SP) that emerge from the long wavelength (LW) bifurcations for $M=2$ and $L=1000\gg\lambda\approx 3.09$ (where $\lambda$ is the wavelength leading to instability of $Q_{2}^{\ast}$). The solution branches are projected with Eq. 25 in (a), the maximum value of $F$ in (b), and the speed $c$ in (c); solid (dashed) lines denote linearly stable (unstable) solutions. The insets in (a) show the locations of several folds along the EP branch that are invisible otherwise, and schematics of the spatial eigenvalues $\zeta$ of $Q_{2}^{\ast}$ at ($s=s_{c}^{t}$) and near (pentagon) the LW onset along the upper SP branch. The square denotes two zero eigenvalues due to mass conservation and spatial reversibility. This also holds for the LW of $Q_{0}^{\ast}$. The right inset in (a) shows the eigenvalue configuration of $Q_{0}^{\ast}$ at a stable EP (i.e., with respective speed $c$, inverted hollow triangle), where the circle shows the zero eigenvalue from mass conservation. The inset in (b) shows the eigenvalue configuration for EPs after the parity-breaking bifurcation. The insets in (c) show solutions at the T-points zoomed into the peaks and the separation of the two T-points. The critical values $(s_{c}^{t},s_{LW},s_{HB},s_{T},s_{P},s_{PD})=(0.409,0.548,0.565,0.601,0.684,0.978)$ denote the codimension-2 LW/finite wavenumber Hopf bifurcation (WB), the LW of $Q_{0}^{\ast}$, the homogeneous Hopf (HB), the T-point bifurcation (red square), the parity-breaking bifurcation of SPs (black triangle), and wavelength-doubling (red circle) bifurcations of EPs. The other symbols (pentagon, diamond, etc.) along the branches represent selected locations of solutions given in Fig. 11. The right semicircle denotes the location of the HB (see Fig. 10). Other parameter values as in Table 1 with $b=b_{c}\approx 0.067$.
@@
Figure 10: (a) Bifurcation diagram computed via numerical continuation as a function of $s$, showing excitable pulses (EPs), homogeneous steady states $Q^{\ast}_{0,1,2}$, and the long wavelength oscillatory (LWO) solutions that emerge from the homogeneous Hopf bifurcation (HB) for $M=2$ and $L=2000$. The solution branches are projected with Equation 25. Solid (dashed) lines denote linearly stable (unstable) solutions. The left inset shows that the propagation speed $c$ of the EPs tends to zero as it approaches the saddle-node $SN_{Q_{1/2}^{\ast}}$. The right insets show schematics of the spatial eigenvalues $\zeta$ at the onset near the saddle-node $SN_{Q_{1/2}^{\ast}}$ and shortly after where $Q_{1}^{\ast}$ is the background state along the crimson red EP branch. The square denotes the two zero eigenvalues due to mass conservation and spatial reversibility when $c=0$ and the circle shows the persistent zero eigenvalue from mass conservation once spatial reversibility is lost. The critical values $(s_{HB},s_{SN}^{1/2},s_{T})=(0.565,0.601,0.601)$ represent the homogeneous Hopf (HB), saddle-node of $Q_{1/2}^{\ast}$, and the T-point, respectively. (b) Real and (c) magnitude of the imaginary parts of the spatial eigenvalues $\zeta$ of the background state $Q_{2}^{\ast}$ along the red branch of EP solutions given between the T-points (TP, square) and the far-left fold (SN, diamond). Here, $\beta$ denotes the arc length along the red EP branch. The sixth eigenvalue is not shown but is negative, with $|\zeta|\sim\mathcal{O}(100)$. The dashed line in (b) denotes the magnitude of the real part of the negative real part complex conjugate pair (orange). The starting TP is given by red square in Fig. 9 and the ending TP by the bordered square. The other symbols (pentagon, diamond, etc.) along the branches represent selected locations of solutions given in Fig. 11d,e. Other parameter values as in Table 1 with $b=b_{c}\approx 0.067$.
@@
Figure 11: Solution profiles at selected locations along the (a) EP, (b) SP, and (c) WP branches in Fig. 9, and the (d) EP and (e) LWO branches in Fig. 10, where for visibility, the whole domain length of $L=1000$ is not always shown. Here $(s_{PD},s_{P},s_{T}{,s_{c}^{t},s_{LW}})\approx(0.978,0.684,0.601{,0.409,0.548})$ are defined as in Fig. 9, $s_{SN}^{1/2}=0.601$ as in Fig. 10, and $(c_{T_{1}},c_{T_{2}})\approx(0.260,0.222)$ denote the propagation speeds at the two T-points. Selected phase space projections are shown for EPs in (f) and WP solutions in (g), where in (g) only the homogeneous steady state $Q_{2}^{\ast}$ exists.
@@
where the full system for which the computation is done, is being described in Equation 42. The inset in Fig. 9b shows the spatial eigenvalue configuration about $Q_{0}^{\ast}$ of the EP branch shortly after the parity-breaking bifurcation but before the onset of saddle-focus connections. Here we see that the zero eigenvalue from spatial reversibility becomes negative but initially all eigenvalues are real. Afterwards, the solutions stabilize after the first fold and destabilize at the subsequent left fold (hollow square). The inset in Fig. 9a shows the spatial eigenvalue configuration about $Q_{0}^{\ast}$ of the EP branch (inverted hollow triangle), where we see that the trailing edge has become oscillatory. After the EPs destabilize, the branch folds again before reaching a T-point bifurcation (square) at $s=s_{T}$ (see Fig. 11a). Here, the solutions form a homoclinic cycle with $Q_{0}^{\ast}$ at the saddle-node, i.e., the solutions briefly plateau at $Q_{2}^{\ast}=Q_{1}^{\ast}$ (see square in Fig. 11a). After passing the T-point, the spatially oscillatory tail becomes a leading edge decaying to $Q_{2}^{\ast}$, as shown by the profile at the left fold (diamond), see also the phase portrait projection in Fig. 11f. This switch is possible by additional annihilation of real eigenvalues and formation of a double focus before reaching the left fold. In Fig. 10b,c, we show the spatial eigenvalues, $\zeta$, corresponding to the above transitions. Here we see that the solutions are saddle-focus connections towards the T-points (squares) and focus-focus connections around the far-left fold (diamond). At the HB of the HSS $Q_{2}^{\ast}$, the magnitudes of the real parts of the leading eigenvalues swap and there is a brief regime where saddle-focus Shil-nikov homoclinics emerge [50]. Then, the branch passes the T-point again, where an additional peak appears out of the leading edge oscillations. The branch associated with a bounded two-peak state goes through two more folds (hollow diamond and circle), after which the peaks start to separate until reaching equidistance, and thus, the branch terminates in a wavelength-doubling bifurcation at $s=s_{PD}$ (a consequence of finite size domain). We omit the EP branch with wavelength $L=500$ as the branch is effectively the same as with wavelength $L=1000$. Note that similar bifurcation structures related to peak doubling near T-point bifurcations have also been observed in an RD model but in the absence of mass conservation [47].
@@
Next, we address the steady solution branch that bifurcates from $s=s_{c}^{t}$. As with high mass ($M=4.5$), the solutions emerge subcritically from $Q_{2}^{\ast}$ and $Q_{0}^{\ast}$ but these onsets can be instabilities when $M=2$, as shown in Figures 4 and 9. From both onsets and with both masses, the bifurcating solutions are spatially localized [9, 69] stationary pulses (SP) that rapidly grow in their amplitudes toward the respective folds [16]. From $Q_{2}^{\ast}$, these localizations are holes while from $Q_{0}^{\ast}$, these are peaks and the onset is close to $SN_{Q^{\ast}_{0/1}}$ (this persists for other values of $b$, see also Fig. 3a). Note that similar behaviour is observed at saddle-nodes in non-gradient systems without mass conservation [18, 70, 98]. In the insets of Fig. 9a, we show that the spatial eigenvalue configuration about $Q_{2}^{\ast}$ and $Q_{0}^{\ast}$ follows the typical configuration of eigenvalues required for localization, whereby a pair of double zero eigenvalues split into two reals, coexisting with an additional positive/negative real pair (i.e., intersection of two two-dimensional manifolds [46]). After the folds, the hole and the peak widen forming WP solutions, where $s$ controls the width. Hence, in between the two folds, the solutions form a homoclinic-like cycle, resembling two locked fronts. However, both plateaus are not HSSs of the system, as shown for mass-conserved systems in [88] and seen by the profile and phase-portrait projection marked by the hollow pentagon in Fig. 11b,g. Stability computations show that the WP solutions are linearly stable from the left fold and up to $s\approx 0.718$, see Figure 9.
@@
Figure 12: Bifurcation diagrams computed via numerical continuation as a function of $s$, showing the homogeneous steady states $Q^{\ast}_{0,1,2}$, and travelling waves (TWs) and wave-pinning (WP) solutions with wavelengths $\lambda$ and $2\lambda$ for $M=2$ and $L=2\lambda\approx 2\cdot 3.09$ (where $\lambda$ is the critical wavelength at the instability onset of $Q_{2}^{\ast}$). The solution branches are projected with the Sobolev norm Eq. 25 in (a) and the propagation speed $c$ in (b); solid (dashed) lines denote linearly stable (unstable) solutions. The notation 2TWλ denotes two spatial copies of travelling waves with wavelength $\lambda$. Here, TWs and WP solutions form a bistable region with wavelengths $\lambda$ and $2\lambda$, respectively. The critical values $(s_{c}^{t},s_{P}^{\lambda},s_{P}^{2\lambda},s_{PD}^{\lambda})=(0.409,0.593,0.664,0.760)$ denote the codimension-2 long wavelength (LW)/finite wavenumber Hopf (WB) onset, the parity-breaking bifurcations (black triangles) at wavelengths $\lambda$ and $2\lambda$, respectively, and the wavelength-doubling (hollow pentagon) bifurcation of $\lambda$. The symbols (circle, diamond, etc.) represent selected locations of solution profiles given in Fig. 13. Other parameter values as in Table 1 with $b=b_{c}\approx 0.067$.
@@
Figure 14: (a) Bifurcation diagram computed via numerical continuation as a function of $s$, showing the homogeneous steady states $Q^{\ast}_{0,1,2}$, and travelling waves (TWs) and wave-pinning (WP) solution branches with wavelengths $3\lambda/2$ and $3\lambda$ for $M=2$ and $L=3\lambda\approx 3\cdot 3.09$ (where $\lambda$ is the critical wavelength at the instability onset of $Q_{2}^{\ast}$). Here we only show projections with the Sobolev norm Eq. 25; solid (dashed) lines denote linearly stable (unstable) solutions. The notation 2TW3λ/2 denotes two spatial copies of TWs with wavelength $3\lambda/2$, where the onsets of each one of the solution branches are at the respective parity-breaking bifurcations. The critical values are $(s_{c}^{t},s_{P_{l}}^{3\lambda/2},s_{P_{r}}^{3\lambda/2},s_{P}^{3\lambda},s_{H}^{3\lambda/2},s_{PD}^{3\lambda/2})=(0.409,0.449,0.629,0.659,0.604,0.821)$, for the codimension-2 long wavelength (LW)/finite wavenumber Hopf (WB) bifurcation, the parity-breaking bifurcation at wavelengths $3\lambda/2$ on the left and right, and $3\lambda$, respectively, and the Hopf (H) and wavelength-doubling bifurcations of TW3λ/2. (b-e) Selected solution profiles for (b) 2WP3λ/2 (c) 2TW3λ/2 (d) WP and (e) TW at corresponding points marked in (a). Other parameter values as in Table 1 with $b=b_{c}\approx 0.067$.